Net Present Value Method, Present Value Index, and Analysis

First United Bank Inc. is evaluating three capital investment projects using the net present value method. Relevant data related to the projects are summarized as follows:

Branch
Office
Expansion
Computer
System
Upgrade
ATM
Kiosk
Expansion
Amount to be invested $754,749 $499,898 $249,409
Annual net cash flows:
Year 1 316,000 224,000 149,000
Year 2 294,000 202,000 103,000
Year 3 269,000 179,000 75,000
Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1. Assuming that the desired rate of return is 10%, prepare a net present value analysis for each proposal. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest dollar.

Branch Office Expansion Computer System Upgrade ATM Kiosk Expansion
Present value of net cash flow total $ $ $
Amount to be invested $ $ $
Net present value $ $ $

2. Determine a present value index for each proposal. If required, round your answers to two decimal places.

Present Value Index
Branch Office Expansion
Computer System Upgrade
Install Internet Bill-Pay

3. Which proposal offers the largest amount of present value per dollar of investment?
Answer #1
1
Branch Office Expansion Computer System Upgrade ATM Kiosk Expansion
Present value of net cash flow total 732107 504897 276844
Amount to be invested 754749 499898 249409
Net present value -22642 4999 27435
2
Present Value Index
Branch Office Expansion 0.97 =732107/754749
Computer System Upgrade 1.01 =504897/499898
ATM Kiosk Expansion 1.11 =276844/249409
3
ATM Kiosk Expansion offers the largest amount of present value per dollar of investment
Workings:
Branch Office Expansion
Cash flows PV factor 10% Present value
1 316000 0.909 287244
2 294000 0.826 242844
3 269000 0.751 202019
Total 732107
Computer System Upgrade
Cash flows PV factor 10% Present value
1 224000 0.909 203616
2 202000 0.826 166852
3 179000 0.751 134429
Total 504897
ATM Kiosk Expansion
Cash flows PV factor 10% Present value
1 149000 0.909 135441
2 103000 0.826 85078
3 75000 0.751 56325
Total 276844